OpenStudy (anonymous):

the function H(x)=sin x cos x/2tan^2x i even. true or false?

OpenStudy (amistre64):

i even? I assume that means "is even"

OpenStudy (amistre64):

even means reflects about the y axis right?

OpenStudy (amistre64):

if f(-x) = f(x) it is an even function x^2 = (-x)^2

OpenStudy (amistre64):

cos is an even function and sin is an odd function in general....just recapping my memory here :)

OpenStudy (anonymous):

yep, that be always true

OpenStudy (amistre64):

sin(x) cos(x) -sin(x) cos(x) ---------- = ----------- ???? 2tan^2(-x) -2tan^2(x)

OpenStudy (amistre64):

i think I messed up that tan^2(x) part lol hold on

OpenStudy (amistre64):

2tan(x)tan(x) = 2-tan(x)-tan(x) = 2tan^2(x) its (+) on the bottom and (-) up top; it aint even then

OpenStudy (anonymous):

"All Students Take Calc" All positive in 1st quad Sine positive in 2nd quad Tangent positive in 3rd quad Cosine positive in 4th quad

OpenStudy (anonymous):

yeah, it's not even it comes out neg

OpenStudy (amistre64):

well of course tan is odd....it looks like x^3 lol right?

OpenStudy (anonymous):

lool so taaan